OpenStudy (mackenzie2013):

What is the radius of a circle that has a circumference of 81.64 meters? Please explain steps!

OpenStudy (anonymous):

Using the formula C=2πr Solving forr r=C 2π=81.64 2·π≈12.99341m

OpenStudy (anonymous):

Sorry Using the formula C=2πr Solving forr r=C/2π=81.64/2·π≈12.99341m